# Comparison of Offset and Ratio Weighted Regressions in Tweedie Models with Application to Mid-Term Cancellations

### v1

@@ 5.2 Offset and Ratio Approaches
To compare the offset and ratio regression approaches, we estimate the parameters using a Tweedie distribution for each segment of the portfolio (insureds with full exposure and those with mid-term cancellation). Furthermore, for each approach (ratio and offset), we calculate the premium for each contract using both methods.
It is worth noting that our dataset is a sample from Boucher and Coulibaly (2024). Accordingly, we applied the same estimated variance parameter, $p=1.42$, as used in their study for premium modelling. As discussed in Section 2, the dispersion parameter does not affect premium modelling; thus, we do not emphasize its role in this analysis.
@@ 5.2.1 Estimated Parameters
The estimated parameters $\beta$ for covariates $X_{1}-X_{8}$ can be different for both approaches. To compare the estimated values, we compute a coefficient ratio for each covariate by dividing the estimated coefficient from the offset approach by the corresponding coefficient from the ratio approach. If the estimated values were identical for both approaches, this ratio would be 1 for all covariates. These results are displayed in Figure 5. Without suprise, the estimated coefficients are identical in both the offset and ratio approaches, which was expected because the offset and ratio approaches yield equivalent results when the risk exposure is equal to 1.
On the other hand, for contracts with mid-term cancellations, the estimated coefficients differ noticeably between the offset and ratio approaches. This divergence is anticipated because, as discussed in Section 3, the log-likelihood and gradient functions differ for the two approaches when risk exposure is less than 1. The significant differences, particularly in coefficients associated with covariates $X1$, $X2$, and $X4$, highlight how the choice of approach impacts premium estimation. These differences imply that premiums for certain policyholder profiles will vary substantially based on the chosen approach.
@@ 5.2.3 Financial Equilibrum Analaysis
In the Full exposure group, the observed gap is identical for both the offset and ratio approaches, explained by the identical estimated premiums in these methods. As previously noted, particularly in Section 4.3.1, it is interesting to observe that the sums of predictions and observations by risk class differ even when risk exposure equals one. In the Mid-term cancellation group, however, some differences emerge in the observed gaps between approaches. Here, the ratio approach appears to perform better, with the sum of estimated premiums aligning more closely with the aggregate observed loss costs than in the offset approach.
Figure 7: Observed gaps comparison (left: Full exposure, right: Mid-term cancellation)
@@ 5.3 Discussion
Even though neither of the two approaches studied is perfect in every respect, the advantages of using the ratio approach for a Tweedie regression are clear for the insurance portfolio we analyzed. However, we also recognize that the analysis of the approach to be taken has brought to light several important issues, which may lead to different choices or potential solutions depending on the context:
  1. 1.
    Approach Choice and Portfolio Composition:
    As shown in Figure 1, the influence of the chosen approach is relatively minor for portfolios with few mid-term cancellations, such as home insurance or certain commercial insurance lines. In these cases, where most contracts have full-year exposures, the offset and ratio approaches produce similar results. However, for insurance portfolios with a higher proportion of contracts having partial-year exposures — typical in certain types of insurance, such as North American car insurance — the choice between the offset and ratio approaches becomes critical.
  2. 2.
    Impact on Premiums and Policyholder Fairness:
    The different coefficient estimates produced by the offset and ratio approaches imply that premiums for the same insured can vary significantly depending on the approach (see the value of estimated parameters from Figure 5). This difference of premium raises important considerations in a competitive insurance market, as it could influence which policyholder profiles benefit or are disadvantaged by each approach. Understanding how these differences impact policyholder fairness and market competitiveness could help insurers decide which premium structure better aligns with market expectations and policyholder satisfaction.
  3. 3.
    Mid-term Cancellations and Loss Costs:
    Table 1 and Figure 3 demonstrate that mid-term cancellations significantly affect loss costs, highlighting the importance of addressing cancellations more precisely in ratemaking models. Rather than simply incorporating risk exposure within a single pricing model, as done with a Tweedie distribution in this paper, analyzing specific causes of cancellations could enable the development of more tailored approaches. By examining the reasons behind cancellations, insurers could create models that account for different cancellation behaviours and adjust premiums accordingly, potentially leading to more accurate pricing strategies.
@@ 6 Conclusion
This study compares the ratio and offset approaches to premium estimation, focusing on how each method incorporates risk exposure. The ratio approach assumes a contract’s premium is directly proportional to its risk exposure, treating the exposure as a weight independent of the response variable’s distribution. In contrast, the offset approach, particularly with a logarithmic link function, incorporates the log of the risk exposure as an offset variable in the premium model. Our findings show that these approaches are not equivalent when the response variable follows a Tweedie distribution.
Through a detailed comparison, we identified the respective strengths and weaknesses of each approach. Additionally, we demonstrated that the offset approach can be reformulated as a weighted regression, structurally similar to the ratio approach. Thus, the distinction between these methods primarily lies in the choice of weights, framing the comparison as one between two weighted regressions.
Our analysis of weight selection revealed that, for any given contract, the weight in the offset approach is consistently greater than that in the ratio approach. This supports our conclusion that the premium estimator from the offset approach generally outperforms the ratio approach, especially when considering bias and variance minimization. However, our assessment of the random gap between aggregate loss costs and the sum of premium estimators suggests that this gap tends to be larger for the offset approach than for the ratio approach.
To evaluate the effectiveness of each approach for pricing, we assessed the balance property, which aligns the sum of estimated premiums with the aggregate actual loss costs. Based on this criterion, the ratio approach performed better than the offset approach. As discussed in Section 5.3, we recommend the ratio approach, particularly in portfolios with diverse risk exposures, due to its ability to better align premiums with actual loss costs. For future research, comparing the premium estimators of both methods using auto-calibration techniques, as suggested by Denuit et al. (2021), Lindholm et al. (2023), Denuit et al. (2024) and Denuit and Trufin (2024), would be valuable, especially given the positive bias observed in both approaches.
Appendix: Descriptive statistics of covariates

### v2

@@ 5.2 Offset and Ratio Approaches
To compare the offset and ratio regression approaches, we estimate the parameters using a Tweedie distribution for each segment of the portfolio (insureds with full exposure and those with mid-term cancellation). Furthermore, for each approach (ratio and offset), we calculate the premium for each contract using both methods. It is worth noting that our dataset is a sample from Boucher and Coulibaly (2024). Accordingly, we applied the same estimated variance parameter, $p=1.42$, as used in their study for premium modelling. As discussed in Section 2, the dispersion parameter does not affect premium modelling; thus, we do not emphasize its role in this analysis.
@@ 5.2.1 Estimated Parameters
The estimated parameters $\bm{\beta}$ for covariates $X_{1}-X_{8}$ can be different for both approaches. To compare the estimated values, we compute a coefficient ratio for each covariate by dividing the estimated coefficient from the offset approach by the corresponding coefficient from the ratio approach. If the estimated values were identical for both approaches, this ratio would be 1 for all covariates. These results are displayed in Figure 7. Without surprise, the estimated coefficients are identical in both the offset and ratio approaches, which was expected because the offset and ratio approaches yield equivalent results when the risk exposure is equal to 1.
On the other hand, for contracts with mid-term cancellations, the estimated coefficients differ noticeably between the offset and ratio approaches. This divergence is anticipated because, as discussed in Section 3, the log-likelihood and gradient functions differ for the two approaches when risk exposure is less than 1. The significant differences, particularly in coefficients associated with covariates $X_{1}$, $X_{2}$, and $X_{4}$, highlight how the choice of approach impacts premium estimation. These differences imply that premiums for certain policyholder profiles will vary substantially based on the chosen approach.
@@ 5.2.3 Financial Equilibrum Analaysis
In the full-exposure group, the sum of individual observed gaps is identical for both the offset and ratio approaches, which is explained by the fact that the estimated premiums are identical under the two methods. As noted previously, it is nevertheless interesting to observe that, in the heterogeneous data, the sums of predicted and observed losses by risk class may differ even when the risk exposure equals one. A similar phenomenon is observed under homogeneous data when the offset approach is used. For the mid-term cancellation group, all ratios are reported in the right panel of Figure 9. The ratio approach yields a closer alignment between the sum of estimated premiums and the aggregate observed loss costs than the offset approach. In particular, the ratio of the sum of estimated premiums to the aggregate observed loss cost under homogeneous data using the offset approach is equal to $1.22$, whereas the ratio approach yields exact aggregate balance under homogeneity. As discussed in Section 4.3.3, ratios of similar magnitude are observed for both approaches in the heterogeneous case.
Figure 9: Sum of individual observed gaps comparison (left: Full exposure, right: Mid-term cancellation)
@@ 5.3 Discussion
Even though neither of the two approaches studied is perfect in every respect, the advantages of using the ratio approach for a Tweedie regression are clear for the insurance portfolio we analyzed. However, the analysis of the modeling choices considered in this paper also brings to light several important issues, which may lead to different conclusions or modeling strategies depending on the portfolio structure and the underlying risk characteristics.
First, the relevance of the offset or ratio approach strongly depends on the composition of the insurance portfolio. As illustrated in Figure 1, the impact of the chosen approach remains relatively limited for portfolios characterized by a low frequency of mid-term cancellations, such as home insurance or certain commercial insurance lines, where most contracts exhibit full-year exposures. In such contexts, the offset and ratio approaches yield similar results. In contrast, for portfolios with a substantial proportion of partial-year exposures—typical of certain insurance markets, such as North American automobile insurance—the choice between the two approaches becomes critical and can materially affect the modeling outcomes.
These differences are not only technical but also have direct implications for premium determination and policyholder fairness. As shown by the estimated parameters reported in Figure 7, the offset and ratio approaches may lead to substantially different premiums for the same insured risk. In a competitive insurance market, such discrepancies may influence which policyholder profiles are favored or disadvantaged by each approach. Understanding how these pricing differences affect fairness perceptions and market competitiveness is therefore an important consideration when selecting a ratemaking methodology.
The empirical results further emphasize the central role of mid-term cancellations in shaping loss costs. Table 1 and Figure 4 show that cancellations have a significant impact on observed loss costs, underlining the need to address this phenomenon more explicitly in pricing models. Rather than incorporating exposure mechanically within a single Tweedie framework, as done in this paper, a more detailed analysis of cancellation behavior could allow insurers to develop more refined modeling strategies. In particular, distinguishing between different causes of cancellations may help capture heterogeneous risk dynamics and improve premium adequacy.
Finally, the analysis raises broader questions regarding the assumption of strict linear proportionality between exposure and expected loss cost, an assumption that underlies both the offset and ratio approaches. While this assumption is widely adopted in actuarial practice, the descriptive results presented in Figure 5 suggest that loss costs do not necessarily scale linearly with exposure. Shorter exposure periods, in particular, appear to be associated with systematically different loss dynamics, consistent with several scenarios discussed in Section 5.1.1. Although both approaches considered here impose linearity by construction, these empirical patterns indicate that relaxing this assumption could provide additional modeling flexibility and lead to a more accurate representation of exposure-related heterogeneity.
@@ 6 Conclusion
This paper has examined two widely used strategies for incorporating risk exposure into premium estimation under Tweedie regression models: the offset approach and the ratio approach. Although both methods rely on the same proportionality assumption between exposure and expected loss, we have shown that they induce fundamentally different weighting structures and therefore lead to distinct statistical and financial properties. From a purely inferential perspective, both approaches yield consistent estimators of the true parameter vector under mild regularity conditions. A first-order asymptotic analysis based on the leading Gaussian term of the Edgeworth expansion suggests that the offset approach may exhibit greater efficiency, in the sense of a smaller asymptotic covariance matrix. Simulation results seem to confirm this comparison, indicating that in finite samples the offset approach also displays smaller empirical variance. Beyond parameter estimation, this paper emphasizes that premium modeling in insurance is inherently a financial exercise, for which aggregate balance considerations play a central role. From this standpoint, a key contribution of this work is to show that the ratio approach exhibits a structurally stronger proximity to financial equilibrium. In homogeneous portfolios, it satisfies an exact aggregate balance condition by construction, while in heterogeneous settings it remains systematically closer to balance than the offset approach. Importantly, this property is not driven by asymptotics or incidental compensation effects, but follows directly from the structure of the estimating equations. In the final section of the paper, we complement the theoretical and simulation-based analyses with an empirical study based on a real Canadian automobile insurance portfolio. This application allows for a direct comparison of the offset and ratio approaches both in terms of parameter estimates—where the two approaches yield different values—and in terms of aggregate financial balance. The empirical results indicate that the ratio approach achieves a closer alignment between the sum of estimated premiums and the aggregate observed loss costs, whereas the offset approach exhibits a systematic imbalance.
The empirical results also suggest that a more refined examination of the relationship between risk exposure and observed loss costs may be warranted. In particular, the assumption of a strictly linear proportionality between exposure and expected loss appears to be restrictive in the presence of mid-term cancellations and heterogeneous exposure patterns, even though such proportionality is implicitly imposed by both modeling approaches considered in this paper. One important source of this heterogeneity arises when vehicles are replaced during the policy term, either following the sale and purchase of another vehicle or, as discussed in Section 5, after a significant loss that renders the original coverage unnecessary. In this context, more detailed tracking of vehicle replacements within insurers’ databases could help improve the representation of exposure-related risk. More generally, these findings indicate that, while risk exposure $t$ remains a vehicle-level quantity, its relationship with risk may depend on the broader exposure configuration of the policy to which the vehicle belongs. In particular, when multiple vehicles are insured sequentially or concurrently under the same policy, considering interactions among vehicle-level exposures may improve the representation of the underlying risk structure.
Appendix 1: Calculation of Asymptotic Covariance Matrices
This appendix summarizes results from White (1982) that are directly relevant to Proposition 2. We provide additional details on the calculation of the asymptotic covariance matrices of the parameter estimators in the Offset and Ratio approaches. To this end, we begin by recalling the expression of the response variance in both approaches, which is fundamental for deriving the covariance matrices:
$\mathrm{Var}\!\left[Z_{i}\mid\mathbf{X}_{i},w_{i}\right]=\frac{\phi}{w_{i}}\zeta_{i}^{p},\quad i=1,\ldots,n,$
where $\zeta_{i}=\exp\{\mathbf{X}_{i}^{\top}\bm{\beta}\}$, and $w_{i}=t_{i}^{2-p}$ under the Offset approach while $w_{i}=t_{i}$ under the ratio approach.
For convenience, we also introduce the notation
$\bm{D}(W)=\operatorname{diag}\!\left(w_{i}\zeta_{i}^{2-p}\right)_{i=1,\ldots,n},$
with the vector $\bm{W}=(w_{1},\ldots,w_{n})^{\top}$. This expression for the variance is used in the computation of the asymptotic covariance matrix of the parameter estimator. Following White (1982), this matrix is defined as
$\Sigma=A^{-1}BA^{-1},$
where $A=-\tfrac{1}{\phi}\bm{X}^{\top}\bm{D}(W)\bm{X}$ and $B=\mathbb{E}\!\left[\nabla(\bm{\beta})\nabla(\bm{\beta})^{\top}\right]$, with the score vector
$\nabla(\bm{\beta})=\frac{1}{\phi}\sum_{i=1}^{n}w_{i}\frac{Z_{i}-\zeta_{i}}{\zeta_{i}^{p-1}}\mathbf{X}_{i}.$
Then, the matrix $B$ is
$B$ $=\frac{1}{\phi^{2}}\mathrm{E}\!\left[\left(\sum_{i=1}^{n}w_{i}\frac{Z_{i}-\zeta_{i}}{\zeta_{i}^{p-1}}x_{i,j_{1}}\right)\left(\sum_{i=1}^{n}w_{i}\frac{Z_{i}-\zeta_{i}}{\zeta_{i}^{p-1}}x_{i,j_{2}}\right)\right]_{j_{1},j_{2}=0,\ldots,q}$
$=\frac{1}{\phi^{2}}\mathrm{E}\!\left[\sum_{i=1}^{n}w_{i}^{2}\frac{(Z_{i}-\zeta_{i})^{2}}{\zeta_{i}^{2p-2}}x_{i,j_{1}}x_{i,j_{2}}\right]_{j_{1},j_{2}=0,\ldots,q}$
$=\frac{1}{\phi^{2}}\sum_{i=1}^{n}w_{i}^{2}\frac{\mathrm{Var}\!\left[Z_{i}\mid\mathbf{X}_{i},w_{i}\right]}{\zeta_{i}^{2p-2}}x_{i,j_{1}}x_{i,j_{2}}$
$=\frac{1}{\phi^{2}}\sum_{i=1}^{n}w_{i}^{2}\frac{\tfrac{\phi}{w_{i}}\zeta_{i}^{p}}{\zeta_{i}^{2p-2}}x_{i,j_{1}}x_{i,j_{2}}$
$=\frac{1}{\phi}\sum_{i=1}^{n}w_{i}\zeta_{i}^{2-p}x_{i,j_{1}}x_{i,j_{2}}$
$=\frac{1}{\phi}\bm{X}^{\top}\bm{D}(W)\bm{X}$
$=-A.$
Hence, B=AB=-A, and the asymptotic covariance matrix reduces to
$\Sigma=-A^{-1}.$